# BODex: Scalable and Efficient Robotic Dexterous Grasp Synthesis Using Bilevel Optimization

### v1

@@ Abstract
Robotic dexterous grasping is a key step toward human-like manipulation. To fully unleash the potential of data-driven models for dexterous grasping, a large-scale, high-quality dataset is essential. While gradient-based optimization offers a promising way for constructing such datasets, existing works suffer from limitations, such as restrictive assumptions in energy design or limited experiments on small object sets. Moreover, the lack of a standard benchmark for comparing synthesis methods and datasets hinders progress in this field. To address these challenges, we develop a highly efficient synthesis system and a comprehensive benchmark with MuJoCo for dexterous grasping. Our system formulates grasp synthesis as a bilevel optimization problem, combining a novel lower-level quadratic programming (QP) with an upper-level gradient descent process. By leveraging recent advances in CUDA-accelerated robotic libraries and GPU-based QP solvers, our system can parallelize thousands of grasps and synthesize over $49$ grasps per second on a single NVIDIA 3090 GPU. Our synthesized grasps for Shadow Hand and Allegro Hand achieve a success rate above $75\%$ in MuJoCo, with a penetration depth and contact distance of under $1$ mm, outperforming existing baselines on nearly all metrics. Compared to the previous large-scale dataset, DexGraspNet, our dataset significantly improves the performance of learning models, with a simulation success rate from around $40\%$ to $80\%$. Real-world testing of the trained model on the Shadow Hand achieves an $81\%$ success rate across 20 diverse objects.
@@ I Introduction
Robotic dexterous grasping is foundational to general manipulation and thus an important research topic. While large-scale data collection and learning-based methods have achieved significant success in parallel gripper grasping [1, 2], their potential for dexterous hands remains largely unexplored, partly due to the absence of a large-scale, high-quality dataset. For parallel grippers, sampling numerous gripper poses and labeling them using analytic metrics is feasible. However, direct sampling for dexterous hands, which often have over 20 degrees of freedom (DoF), is impractical. Furthermore, analytic metrics for dexterous hands are less effective due to the complex contact interactions with objects.
To address these challenges and construct a grasp dataset for dexterous hands, recent studies have explored gradient-based optimization combined with simulation validation as a promising approach. However, existing optimization methods face several limitations. Some works [3, 4] rely on restrictive assumptions in the energy, such as equal contact forces and no friction, while others [5, 6] study on a limited set of objects. Moreover, differences in robot hands, simulators, and evaluation metrics across studies make comparison difficult.
This work aims to overcome these challenges by developing an efficient grasp synthesis system and a comprehensive benchmark. We formulate grasp synthesis as a bilevel optimization problem: the lower-level quadratic programming (QP) determines the optimal force combination for each contact at the current hand pose to achieve a desired wrench, while the upper-level process performs gradient descent on the hand pose to minimize the difference between the desired wrench and the best-applied wrench, as determined by the lower-level QP. Our system also supports synthesizing pre-grasp poses that maintain a certain distance from the object, aiding in planning collision-free hand-arm trajectories and controlling the hand to apply force on the object.
@@
Dataset Hand Sim./Real Objects Grasps Table Pre-grasps Collision-free hand-arm trajectory Method
@@
BODex-Floating Shadow Sim. 2440 3.62M Optimization
BODex-Tabletop Shadow Sim. 2440 3.41M 2.62M Optimization
@@
Finally, we establish a benchmark with the MuJoCo simulator [14] to compare various analytic-based synthesis pipelines, grasp energy functions, and learning-based methods. As shown in Fig. 1, our system significantly outperforms previous baselines across nearly all metrics, achieving a 50x speedup in synthesizing higher-quality grasps. Compared to previous grasp energies, our QP-based energy does not rely on assumptions about contact forces and friction, resulting in a higher success rate and better correlation with simulation outcomes (i.e. a grasp with lower energy is more likely to succeed in simulation). The learning-based method trained on our dataset also greatly outperforms the same model trained on the previous large-scale dataset, DexGraspNet, improving success rates from around $40\%$ to $80\%$. Real-world testing of the trained model on the Shadow Hand achieves an $81\%$ success rate across $20$ diverse objects.
@@ II Related Work
Bilevel optimization for grasp synthesis [5, 20] involves a lower-level optimization process and an upper-level optimization that depends on the outcome of the lower-level process. Therefore, it typically has higher computational costs than methods whose energies do not require optimization. To address this, we implement a GPU-based QP solver [8] capable of solving lower-level QPs in parallel, preventing it from becoming a speed bottleneck.
@@ IV-A Bilevel Optimization Formulation for Grasp Synthesis
By finding the optimal contact forces $\mathbf{y}_{j}$ for a desired wrench $\beta\mathbf{t}_{j}$, the grasp energy $Q_{j}(\mathbf{x})$ measures the difference between the desired wrench and the best-applied wrench $\Sigma_{i=1}^{m}\mathbf{G}_{i}\mathbf{f}_{j,i}$ at the hand pose $\mathbf{x}$. To get a better hand pose, the upper-level process performs gradient descent on $Q_{j}$, which is differentiable to $\mathbf{x}$ because $\mathbf{G}_{i}$ is differentiable to $\mathbf{x}$. The final grasp energy sums up several $Q_{j}$ with different $\mathbf{t}_{j}$ to encourage the applicable wrenches in multiple directions. For a force-closure grasp, $\{\mathbf{t}_{j}\}$ is set to be the six unit vectors along the positive and negative 3D force axes, with zero 3D torques, e.g., $[1,0,0,0,0,0]$ and $[-1,0,0,0,0,0]$. If a grasp can resist these six $\mathbf{t}_{j}$, it can resist the object gravity in any direction by a linear combination of these $\mathbf{t}_{j}$. Our formulation is also more flexible in customizing desired wrenches than the predefined primitive in TDG [19].
@@ IV-B Solving the Bilevel Optimization for Grasp Synthesis
In each iteration of the upper-level optimization, the hand pose $\mathbf{x}$ is used to compute the transformation of each hand link, represented as $\mathbf{R}_{i}$ and $\mathbf{T}_{i}$, by forward kinematics (FK). This yields the expected hand contact points in the world frame, given by $\mathbf{c}_{i,w}=\mathbf{R}_{i}\mathbf{c}_{i,l}+\mathbf{T}_{i}$. Next, the object points $\mathbf{p}_{i}$ and normals $\mathbf{n}_{i}$ are calculated by nearest-point query, as shown in the left of Fig. 2. They are then used to construct the grasp matrix $\mathbf{G}_{i}$ in Eq. 2 and the QPs in Eq. 7.
@@ IV-C Coarse-to-fine Contact Modeling
To accelerate computation, we use HPP-FCL [29], a state-of-the-art library implementing the GJK algorithm, along with OpenMP for parallelization. Since the GJK algorithm is limited to convex meshes, the object is decomposed into multiple convex parts. To reduce unnecessary computations, we introduce a broad-phase step that calculates the distance between the oriented bounding box (OBB) of each object part and the bounding sphere of the fingertip collision mesh. Only object parts with OBB-to-sphere distances less than the distance from the fingertip’s sphere approximation to the entire object are considered for the GJK algorithm.
@@ V-A Simulation Environment and Object Assets
Object assets are taken from DexGraspNet [4]. All objects are preprocessed using ManifoldPlus [33] and normalized so that the diagonal of their bounding boxes measures $2m$. To address the issue of objects being too flat to grasp on a table, we filtered out those with a shortest bounding box edge less than $0.2m$, resulting in $2,440$ valid objects. Each object is then rescaled to four different sizes: $0.06$, $0.08$, $0.10$, and $0.12$, resulting in a total of $9,760$ scaled objects for grasping.
@@ V-B Evaluation Metrics
Speed (S) (unit: $grasp/s$) measures the number of grasps synthesized per second. Our speed is tested on an NVIDIA GeForce RTX 3090 GPU with Intel Xeon Platinum 8375C CPUs, while the speed of baselines came from their papers.
Penetration Depth (PD) (unit: $mm$) measures the maximum intersection distance between the object and the hand. The GJK algorithm is used for the following three metrics. The object is decomposed into convex parts by CoACD [34] while the hand’s collision mesh is convex by design.
@@
Contact Distance (CD) (unit: $mm$) shows the proximity of fingers to the object, calculated as the average minimum distance between the object and each finger’s collision mesh. Fingers that intersect the object are assigned a distance of 00.
@@ V-C1 Comparison with previous pipelines
We compare with DexGraspNet (DGN) [4], SpringGrasp [6], and FRoGGeR [5], using the 16-DOF 4-fingered Allegro Hand, the only hand type supported by the open-source codes of all baselines. For each scaled object in a floating state without a table, $10$ grasps are synthesized, resulting in $97,600$ grasps per method. All experiments in Sec. V-C follows this setting.
For baselines without pre-grasp and squeezed poses, we generate these poses by scaling the grasp poses by factors of $0.9$ and $1.1$, respectively. This does not greatly impact the performance of the Allegro Hand, as shown in Tab. III.
Since each baseline is designed and evaluated using its own criteria, achieving a fully fair comparison, particularly for simulation success rate, is challenging. To mitigate this, we relax the rotation threshold for all baselines to $180^{\circ}$.
Despite this concession, our pipeline still greatly outperforms previous methods across nearly all metrics, especially in simulation success rate and speed, as shown in Table II and Fig. 1. Moreover, our grasps show lower penetration depth and contact distances, while maintaining comparable diversity. Visualizations are provided in Fig. 3.
Name SSR $\uparrow$ $\uparrow$ PD $\downarrow$ CD $\downarrow$ SPD $\downarrow$ FVR $\downarrow$
DexGraspNet 31.37 0.40 13.5 6.90 0.93 0.31
SpringGrasp 37.24 0.67 16.2 6.18 0.48 0.44
FRoGGeR 41.97 1.20 2.17 0.88 0.0002 0.28
Ours 89.55 61.5 0.37 0.28 0.008 0.32
@@ V-C2 Comparison with previous grasp energies
For evaluation, we assess all grasps synthesized by the above four energies. A good energy should ensure that grasps with lower energy are more likely to succeed in simulation. This is measured by the ROC curve, which plots the true positive rate versus the false positive rate at varying energy thresholds. A larger area under the ROC curve indicates better performance as a grasp evaluation metric.
The results, shown in Fig. 5, demonstrate that our energy greatly outperforms DFC and TDG for both synthesis and evaluation, as our energy does not assume equal contact forces. Compared to QP_baseline, our energy improves SSR by $10\%$ during synthesis, while the evaluation performance is comparable. Interestingly, as the object scale increases, the success rate decreases. This occurs because smaller objects are easier for the hand to form a wrapping grasp and achieve force closure, whereas larger objects make optimization harder. Notably, this decline is less pronounced with our energy, highlighting its advantages. Moreover, this phenomenon explains why the Allegro Hand achieves a higher success rate than the Shadow Hand, as it is considerably larger.
Hand Coarse-to-fine Pre-grasp SSR S PD CD
Shadow 57.8 52.7 0.59 0.42
Shadow 68.4 55.3 0.64 0.87
Shadow 75.8 49.8 0.16 0.27
Allegro 83.2 65.2 0.38 0.68
Allegro 86.7 67.4 0.66 1.01
Allegro 89.6 61.5 0.37 0.28
@@ V-C3 Ablation Study
As shown in Table III, the pre-grasp and coarse-to-fine strategies improve the simulation success rate, penetration depth, and contact distance, at the cost of speed. The simulation success rate of the Allegro Hand is less affected by these changes, likely due to the robustness of its spherical end effector. Moreover, the Allegro Hand is faster than Shadow Hand due to fewer fingers. Notably, all hyperparameters are kept the same for these two hands.
Dataset Method SSR $\uparrow$ PD $\downarrow$ CD $\downarrow$ SPD $\downarrow$ FVR $\downarrow$
Ours-F GraspTTA 14.81 27.27 8.89 0.05 0.39
Ours-F ISAGrasp 20.74 17.31 5.69 0.006 0.38
Ours-F DP3 61.06 6.17 2.55 0.04 0.50
Ours-F UDG 80.12 4.31 1.64 0.001 0.38
DGN UDG 42.02 8.50 1.12 0.10 0.34
UDG UDG 36.85 12.03 1.19 0.16 0.32
Ours-T UDG 78.80 6.44 0.98 0.02 0.33
@@ V-D Benchmarking Learning-based Grasp Synthesis
We benchmark $4$ supervised learning architectures: ISAGrasp [22], GraspTTA [21], 3D Diffusion policy [35], and UnidexGrasp (UDG) [23]. They are representative due to their diverse architectures, ranging from naive regression and CVAE to diffusion models and normalizing flows.
To ensure a fair comparison, we standardize the backbone across these methods by a 3D SparseConv network [36] to extract global features from the input partial point cloud. Another change is that the network learns to predict both pre-grasp and grasp pose. We also simplify the complex pipeline of UniDexGrasp and only use the GraspGlow module combined with Mobius Flow [37] for orientation generation. For diffusion models and normalizing flows, we select the top 10 grasps from 100 samples for testing, as these models allow for probability estimation of each sample, which indicates the quality of the prediction, unlike the other two methods.
The $2,440$ objects are randomly split into a training set and a testing set in a 4:1 ratio. $4$ scales are applied to each object, and $5$ partial point clouds are rendered for each scaled object from random camera viewpoints. Each model is trained with $50,000$ iterations with a batch size of $256$.

### v2

@@ Abstract
Robotic dexterous grasping is important for interacting with the environment. To unleash the potential of data-driven models for dexterous grasping, a large-scale, high-quality dataset is essential. While gradient-based optimization offers a promising way for constructing such datasets, previous works suffer from limitations, such as inefficiency, strong assumptions in the grasp quality energy, or limited object sets for experiments. Moreover, the lack of a standard benchmark for comparing different methods and datasets hinders progress in this field. To address these challenges, we develop a highly efficient synthesis system and a comprehensive benchmark with MuJoCo for dexterous grasping. We formulate grasp synthesis as a bilevel optimization problem, combining a novel lower-level quadratic programming (QP) with an upper-level gradient descent process. By leveraging recent advances in CUDA-accelerated robotic libraries and GPU-based QP solvers, our system can parallelize thousands of grasps and synthesize over $49$ grasps per second on a single 3090 GPU. Our synthesized grasps for Shadow, Allegro, and Leap hands all achieve a success rate above $75\%$ in simulation, with a penetration depth under $1$ mm, outperforming existing baselines on nearly all metrics. Compared to the previous large-scale dataset, DexGraspNet, our dataset significantly improves the performance of learning models, with a success rate from around $40\%$ to $80\%$ in simulation. Real-world testing of the trained model on the Shadow Hand achieves an $81\%$ success rate across 20 diverse objects. The codes and datasets are released on our project page: https://pku-epic.github.io/BODex.
@@ I Introduction
Robotic dexterous grasping is foundational to interacting with the environment and thus an important research topic. While large-scale data collection and learning-based methods have achieved significant success in parallel gripper grasping [1, 2], their potential for dexterous hands remains largely unexplored, partly due to the increased difficulty of data collection. Unlike parallel grippers, dexterous hands often have over 20 degrees of freedom (DoF), thus greatly reducing the effectiveness of directly sampling grasp poses.
Although gradient-based optimization has been explored recently as a promising approach to scaling up the grasp data for dexterous hands, previous methods face several limitations. Some [3, 4] rely on strong assumptions in the grasp quality energy, such as equal contact forces and no friction, while others [5, 6] only study on a limited set of objects. Moreover, the synthesis speed of previous works is quite slow. In addition, differences in robot hands, simulators, and evaluation metrics across studies make comparison difficult.
To address these challenges, we develop an efficient grasp synthesis system and a comprehensive benchmark. We formulate grasp synthesis as a bilevel optimization problem: the lower-level quadratic programming (QP) determines the optimal force combination for each contact at the current hand pose to achieve a desired wrench, without any assumption, while the upper-level process performs gradient descent on the hand pose to minimize the difference between the desired wrench and the best-applied wrench, as determined by the lower-level QP. Our system can also synthesize pre-grasp poses that maintain a certain distance from the object, aiding in planning collision-free hand-arm trajectories and controlling the hand to apply force on the object.
@@
Dataset Hand Sim./Real Objects Grasps Table Pre-grasps Collision-free trajectory Method
@@
BODex-Floating Shadow, Allegro, Leap Sim. 2397 4.09M Optimization
BODex-Tabletop UR10e+Shadow Sim. 2397 3.08M 3.08M Optimization
@@
Finally, we establish a benchmark with MuJoCo [14] to compare various analytic-based grasp synthesis pipelines, grasp energy functions, and learning-based methods. As shown in Fig. 1, our system significantly outperforms baselines across nearly all metrics, achieving a 50x speedup in synthesizing higher-quality grasps. Compared to previous grasp energies, our QP-based energy does not rely on assumptions about contact forces and friction, resulting in a higher success rate and better correlation with simulation outcomes (i.e. a grasp with lower energy is more likely to succeed in simulation). The learning-based method trained on our dataset also greatly outperforms the same model trained on the previous large-scale dataset, DexGraspNet, improving success rates from around $40\%$ to $80\%$. Real-world testing of the trained model on the Shadow Hand achieves an $81\%$ success rate across $20$ diverse objects.
@@ II Related Work
Bilevel optimization for grasp synthesis [5, 20] involves an upper-level optimization process that solves a lower-level optimization problem at each iteration. This formulation nests two optimization problems and differs from sequential optimization [28]. It typically incurs higher computational costs than methods whose grasp energies do not require optimization. To address this, we implement a GPU-based QP solver [8] capable of solving lower-level QPs in parallel, preventing it from becoming a speed bottleneck.
@@ IV-A Bilevel Optimization Formulation for Grasp Synthesis
By finding the optimal contact forces $\mathbf{y}_{j}$ for a desired wrench $\beta\mathbf{t}_{j}$, the grasp energy $Q_{j}(\mathbf{x})$ measures the difference between the desired wrench and the best-applied wrench $\Sigma_{i=1}^{m}\mathbf{G}_{i}\mathbf{f}_{j,i}$ at the grasp pose $\mathbf{x}$. To get a better grasp pose, the upper-level process performs gradient descent on $Q_{j}$, which is differentiable to $\mathbf{x}$ because $\mathbf{G}_{i}$ is differentiable to $\mathbf{x}$. The final grasp energy sums up several $Q_{j}$ with different $\mathbf{t}_{j}$ to encourage the applicable wrenches in multiple directions. For a force-closure grasp, $\{\mathbf{t}_{j}\}$ is set to be the six unit vectors along the positive and negative 3D force axes, with zero 3D torques, e.g., $[1,0,0,0,0,0]$ and $[-1,0,0,0,0,0]$. If a grasp can resist these six $\mathbf{t}_{j}$, it can resist the object gravity in any direction by a linear combination of these $\mathbf{t}_{j}$. Our formulation is also more flexible in customizing desired wrenches than the predefined primitive in TDG [19].
@@ IV-B Solving the Bilevel Optimization for Grasp Synthesis
In each iteration of the upper-level optimization, the grasp pose $\mathbf{x}$ is used to compute the transformation of each hand link, represented as $\mathbf{R}_{i}$ and $\mathbf{T}_{i}$, by forward kinematics (FK). This yields the expected hand contact points in the world frame, given by $\mathbf{c}_{i,w}=\mathbf{R}_{i}\mathbf{c}_{i,l}+\mathbf{T}_{i}$. Next, the object points $\mathbf{p}_{i}$ and normals $\mathbf{n}_{i}$ are calculated by nearest-point query, as shown in the left of Fig. 2. They are then used to construct the grasp matrix $\mathbf{G}_{i}$ in Eq. 2 and the QPs in Eq. 7.
@@ IV-C Coarse-to-fine Contact Modeling
To accelerate computation, we use Coal [30], a state-of-the-art library implementing the GJK algorithm, along with OpenMP for parallelization. Since the GJK algorithm is limited to convex meshes, the object is decomposed into multiple convex parts. To reduce unnecessary computations, we introduce a broad-phase step that calculates the distance between the oriented bounding box (OBB) of each object part and the bounding sphere of the fingertip collision mesh. Only object parts with OBB-to-sphere distances less than the distance from the fingertip’s sphere approximation to the entire object are considered for the GJK algorithm.
@@ V-A Simulation Environment and Object Assets
Object assets are taken from DexGraspNet [4]. All objects are pre-processed and normalized so that the diagonal of their bounding boxes measures $2m$. To address the issue of objects being too flat to grasp on a table, we filtered out those with a shortest bounding box edge less than $0.2m$, resulting in $2,397$ valid objects. Each object is then rescaled to four different sizes: $0.06$, $0.08$, $0.10$, and $0.12$, resulting in a total of $9,588$ scaled objects for grasping.
@@ V-B Evaluation Metrics
Speed (S) (unit: $grasp/s$) measures the number of grasps synthesized per second. Our speed is tested on an NVIDIA GeForce RTX 3090 GPU with Intel Xeon Platinum 8255C CPUs, while the speed of baselines came from their papers.
Penetration Depth (PD) (unit: $mm$) measures the maximum intersection distance between the object and the hand, calculated in MuJoCo.
@@
Contact Distance Consistency (CDC) (unit: $mm$) measures the delta between the maximum and minimum signed distances across all fingers. This metric quantifies the variation in contact distance across different fingers and is invariant to penetration.
@@ V-C1 Comparison with previous pipelines
We compare with DexGraspNet (DGN) [4], SpringGrasp [6], and FRoGGeR [5], using the 16-DOF 4-fingered Allegro Hand, the only hand type supported by the open-source codes of all baselines. For each scaled object in a floating state without a table, $10$ grasps are synthesized, resulting in $95,880$ grasps per method. For baselines without pre-grasp and squeezed poses, we generate these poses by scaling the grasp poses by factors of $0.9$ and $1.1$, respectively. All experiments in this section follow this setting.
As shown in Table II and Fig. 1, our pipeline greatly outperforms previous methods across nearly all metrics, especially in simulation success rate and speed. Moreover, our grasps show lower penetration depth and contact distance consistency, indicating a better contact convergence. Our diversity is also comparable to previous works. Visual comparisons are provided in Fig. 3.
Name SSR $\uparrow$ $\uparrow$ PD $\downarrow$ CDC $\downarrow$ SPD $\downarrow$ FVR $\downarrow$
DexGraspNet 37.0 0.40 4.49 5.53 1.30 32.0
SpringGrasp 12.1 0.67 11.7 20.1 0.80 44.5
FRoGGeR 17.2 1.20 2.14 3.96 0.01 28.9
Ours 79.9 61.5 0.51 2.98 0.01 33.2
@@ V-C2 Comparison with previous grasp energies
For evaluation, we assess all grasps synthesized by the above four energies. A good grasp quality energy should ensure that grasps with lower energy are more likely to succeed in simulation. This is measured by the ROC curve, which plots the true positive rate versus the false positive rate at varying energy thresholds. The grasp energy with a larger area under the ROC curve (AUC) is better.
The results, shown in Fig. 5, demonstrate that our energy greatly outperforms DFC and TDG for both synthesis and evaluation, as our energy does not assume equal contact forces. Compared to QP_baseline, our energy improves SSR by $10\%$ during synthesis, while the evaluation performance is comparable. Interestingly, as the object scale increases, the success rate decreases. This occurs because smaller objects are easier for the hand to form a wrapping grasp and achieve force closure, whereas larger objects make optimization harder. Notably, this decline is less pronounced with our energy, highlighting its advantages.
Hand Coarse2fine Pre-grasp SSR S PD CDC
Shadow 57.8 52.7 0.69 3.07
Shadow 68.4 55.3 0.81 3.62
Shadow 75.8 49.8 0.66 2.83
Allegro 68.2 65.2 0.57 3.16
Allegro 78.1 67.4 0.63 3.42
Allegro 79.9 61.5 0.51 2.98
@@ V-C3 Ablation Study
As shown in Table III, the pre-grasp and coarse-to-fine strategies improve the simulation success rate, penetration depth, and contact distance consistency, at the cost of speed. Moreover, the Allegro Hand is faster than Shadow Hand due to fewer fingers. Notably, all hyperparameters are kept the same for these two hands. We have also synthesized the grasps for Leap hand and released them on our project page, whose quantitative numbers are similar and thus not included here.
Dataset Method SSR $\uparrow$ PD $\downarrow$ CDC $\downarrow$ SPD $\downarrow$ FVR $\downarrow$
Ours-F GTTA 14.8 27.3 29.6 0.15 39.2
Ours-F ISAG 20.7 17.3 22.4 0.08 38.7
Ours-F DP3 62.4 8.91 13.9 0.10 32.6
Ours-F UDG 80.1 3.47 8.85 0.04 33.4
DGN UDG 42.6 8.69 12.5 0.13 33.2
UDG UDG 36.8 12.0 16.5 0.16 32.7
Ours-T UDG 78.8 5.79 9.79 0.05 33.9
@@ V-D Benchmarking Learning-based Grasp Synthesis
We benchmark $4$ supervised learning architectures: ISAGrasp (ISAG) [22], GraspTTA (GTTA) [21], 3D Diffusion policy (DP3) [33], and UnidexGrasp (UDG) [23]. They are representative due to their diverse architectures, ranging from naive regression and CVAE to diffusion models and normalizing flows.
To ensure a fair comparison, we standardize the backbone across these methods by a 3D SparseConv network [34] to extract global features from the input single-view partial point cloud. Another change is that the network learns to predict both pre-grasp and grasp pose. We also simplify the complex pipeline of UniDexGrasp and only use the GraspGlow module combined with Mobius Flow [35] for orientation generation. For diffusion models and normalizing flows, we select the top 10 grasps from 100 samples for testing, as these models allow for probability estimation of each sample, which indicates the quality of the prediction.
The objects are randomly split into a training set and a testing set in a 4:1 ratio. $4$ scales are applied to each object, and $5$ partial point clouds are rendered for each scaled object from random camera viewpoints. Each model is trained with $50,000$ iterations with a batch size of $256$.